It is approximately 480 miles from Los Angeles, California, to San Francisco, California. Allowing for various traffic conditions, a driver can average approximately 60 miles per hour.
a)Write a linear function that expresses the distance traveled, d, as a function of time, t.
b)How far have you traveled after 4 hours?
c)Write a linear function that expresses the distance to be traveled to reach San Francisco, s, as a function of time, t.
d)How far will I need to travel to reach San Francisco after I have traveled 3 hours?

Answer by venugopalramana(3286)

relation between speed(s) ,time(t) and distance(d) is
d=s*t...since s=60 mph...d=60t......
d=480 miles..and s=60.mph....hence
480=s*t..........and........480=60*t
after 4 hrs d=60*4=240 miles
after 3 hrs d=60*3=180 miles
total distance =480 miles.
so remaining distance =480-180=300miles